# FlexGuard: A Design Space for On-Body Feedback for Injury Prevention in Strength Training

### v1

@@ Acknowledgements.
We thank all of the trainers and trainees who participated in our study and shared their time and expertise.References
@@
- Amershi et al. (2014) Saleema Amershi, Maya Cakmak, W. Bradley Knox, and Todd Kulesza. 2014. Power to the People: The Role of Humans in Interactive Machine Learning. AI Magazine 35, 4 (2014), 105–120. doi:10.1609/aimag.v35i4.2513 arXiv:https://onlinelibrary.wiley.com/doi/pdf/10.1609/aimag.v35i4.2513
@@
- Braun and Clarke (2019) Virginia Braun and Victoria Clarke. 2019. Reflecting on reflexive thematic analysis. Qualitative Research in Sport, Exercise and Health 11, 4 (2019), 589–597. doi:10.1080/2159676X.2019.1628806 arXiv:https://doi.org/10.1080/2159676X.2019.1628806
@@
- Li et al. (2017) Shuguang Li, Daniel M. Vogt, Daniela Rus, and Robert J. Wood. 2017. Fluid-driven origami-inspired artificial muscles. Proceedings of the National Academy of Sciences 114, 50 (2017), 13132–13137. doi:10.1073/pnas.1713450114 arXiv:https://www.pnas.org/doi/pdf/10.1073/pnas.1713450114
- Li and Luo (2024) Yunzhu Li and Yiyue Luo. 2024. Intelligent textiles are looking bright. Science 384, 6691 (2024), 29–30. doi:10.1126/science.ado5922 arXiv:https://www.science.org/doi/pdf/10.1126/science.ado5922
@@
- Wang and Li (2023) Jun Wang and Suyi Li. 2023. Building intelligence in the mechanical domain—Harvesting the reservoir computing power in origami to achieve information perception tasks. Advanced Intelligent Systems 5, 9 (2023), 2300086.
@@
- Young et al. (2019) Eric M. Young, Amirhossein H. Memar, Priyanshu Agarwal, and Nick Colonnese. 2019. Bellowband: A Pneumatic Wristband for Delivering Local Pressure and Vibration. In 2019 IEEE World Haptics Conference (WHC). 55–60. doi:10.1109/WHC.2019.8816075

### v2

@@ Acknowledgements.
This work was supported by the Center for Human-Computer Interaction at Virginia Tech (CHCI@VT). We thank our participants for their valuable insights. We also thank Utkarsha Mohan, Xiaohang Tang, and Tong Wu for generously lending study equipment.References
@@
- Al-Sada et al. (2020) Mohammed Al-Sada, Keren Jiang, Shubhankar Ranade, Mohammed Kalkattawi, and Tatsuo Nakajima. 2020. HapticSnakes: multi-haptic feedback wearable robots for immersive virtual reality. Virtual Real. 24, 2 (June 2020), 191–209. doi:10.1007/s10055-019-00404-x
@@
- Amershi et al. (2014) Saleema Amershi, Maya Cakmak, W. Bradley Knox, and Todd Kulesza. 2014. Power to the People: The Role of Humans in Interactive Machine Learning. AI Magazine 35, 4 (2014), 105–120. arXiv:https://onlinelibrary.wiley.com/doi/pdf/10.1609/aimag.v35i4.2513 doi:10.1609/aimag.v35i4.2513
@@
- Braun and Clarke (2019) Virginia Braun and Victoria Clarke. 2019. Reflecting on reflexive thematic analysis. Qualitative Research in Sport, Exercise and Health 11, 4 (2019), 589–597. arXiv:https://doi.org/10.1080/2159676X.2019.1628806 doi:10.1080/2159676X.2019.1628806
@@
- Choi et al. (2022) Kyung Yun Choi, Neska ElHaouij, Jinmo Lee, Rosalind W. Picard, and Hiroshi Ishii. 2022. Design and Evaluation of a Clippable and Personalizable Pneumatic-haptic Feedback Device for Breathing Guidance. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 6, 1, Article 6 (March 2022), 36 pages. doi:10.1145/3517234
@@
- Dill et al. (2023) Sebastian Dill, Andreas Rösch, Maurice Rohr, Gökhan Güney, Luisa De Witte, Elias Schwartz, and Christoph Hoog Antink. 2023. Accuracy evaluation of 3D pose estimation with mediapipe pose for physical exercises. In Current Directions in Biomedical Engineering, Vol. 9. De Gruyter, 563–566.
- Elsayed et al. (2020) Hesham Elsayed, Martin Weigel, Florian Müller, Martin Schmitz, Karola Marky, Sebastian Günther, Jan Riemann, and Max Mühlhäuser. 2020. VibroMap: Understanding the Spacing of Vibrotactile Actuators across the Body. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 4, 4, Article 125 (Dec. 2020), 16 pages. doi:10.1145/3432189
@@
- Hassan et al. (2017) Mahmoud Hassan, Florian Daiber, Frederik Wiehr, Felix Kosmalla, and Antonio Krüger. 2017. FootStriker: An EMS-based Foot Strike Assistant for Running. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 1, 1, Article 2 (March 2017), 18 pages. doi:10.1145/3053332
@@
- Höök et al. (2017) Kristina Höök, Caroline Hummels, Katherine Isbister, Patrizia Marti, Elena Márquez Segura, Martin Jonsson, Florian ’Floyd’ Mueller, Pedro A.N. Sanches, Thecla Schiphorst, Anna Ståhl, Dag Svanaes, Ambra Trotto, Marianne Graves Petersen, and Youn-kyung Lim. 2017. Soma-Based Design Theory. In Proceedings of the 2017 CHI Conference Extended Abstracts on Human Factors in Computing Systems (Denver, Colorado, USA) (CHI EA ’17). Association for Computing Machinery, New York, NY, USA, 550–557. doi:10.1145/3027063.3027082
@@
- Kim et al. (2023) Jiha Kim, Younho Nam, Jungeun Lee, Young-Joo Suh, and Inseok Hwang. 2023. ProxiFit: Proximity Magnetic Sensing Using a Single Commodity Mobile toward Holistic Weight Exercise Monitoring. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 7, 3, Article 105 (Sept. 2023), 32 pages. doi:10.1145/3610920
@@
- Lee et al. (2026) Hsuanling Lee, Jiakun Yu, Shurui Zheng, Te-Yan Wu, and Liang He. 2026. FluxLab: Creating 3D Printable Shape-Changing Devices with Integrated Deformation Sensing. In Proceedings of the Twentieth International Conference on Tangible, Embedded, and Embodied Interaction (Chicago, IL, USA) (TEI ’26). Association for Computing Machinery, New York, NY, USA, 1–12. doi:10.1145/3731459.3773331
@@
- Ley-Flores et al. (2024) Judith Ley-Flores, Laia Turmo Vidal, Elena Márquez Segura, Aneesha Singh, Frederic Bevilacqua, Francisco Cuadrado, Joaquín Roberto Díaz Durán, Omar Valdiviezo-Hernández, Milagrosa Sánchez-Martin, and Ana Tajadura-Jiménez. 2024. Co-Designing Sensory Feedback for Wearables to Support Physical Activity through Body Sensations. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 8, 1, Article 40 (March 2024), 31 pages. doi:10.1145/3643499
- Li et al. (2017) Shuguang Li, Daniel M. Vogt, Daniela Rus, and Robert J. Wood. 2017. Fluid-driven origami-inspired artificial muscles. Proceedings of the National Academy of Sciences 114, 50 (2017), 13132–13137. arXiv:https://www.pnas.org/doi/pdf/10.1073/pnas.1713450114 doi:10.1073/pnas.1713450114
- Li and Luo (2024) Yunzhu Li and Yiyue Luo. 2024. Intelligent textiles are looking bright. Science 384, 6691 (2024), 29–30. arXiv:https://www.science.org/doi/pdf/10.1126/science.ado5922 doi:10.1126/science.ado5922
@@
- Sung et al. (2024) Youjin Sung, Rachel Kim, Kun Woo Song, Yitian Shao, and Sang Ho Yoon. 2024. HapticPilot: Authoring In-situ Hand Posture-Adaptive Vibrotactile Feedback for Virtual Reality. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 7, 4, Article 179 (Jan. 2024), 28 pages. doi:10.1145/3631453
@@
- Wang et al. (2020) Chi Wang, Da-Yuan Huang, Shuo-Wen Hsu, Cheng-Lung Lin, Yeu-Luen Chiu, Chu-En Hou, and Bing-Yu Chen. 2020. Gaiters: Exploring Skin Stretch Feedback on Legs for Enhancing Virtual Reality Experiences. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–14. doi:10.1145/3313831.3376865
- Wang et al. (2023) Guanyun Wang, Yue Yang, Mengyan Guo, Kuangqi Zhu, Zihan Yan, Qiang Cui, Zihong Zhou, Junzhe Ji, Jiaji Li, Danli Luo, Deying Pan, Yitao Fan, Teng Han, Ye Tao, and Lingyun Sun. 2023. ThermoFit: Thermoforming Smart Orthoses via Metamaterial Structures for Body-Fitting and Component-Adjusting. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 7, 1, Article 31 (March 2023), 27 pages. doi:10.1145/3580806
@@
- Wang et al. (2025) Zeyu Wang, Ruotong Yu, Xiangyang Wang, Jiexin Ding, Jiankai Tang, Jun Fang, Zhe He, Zhuojun Li, Tobias Röddiger, Weiye Xu, Xiyuxing Zhang, Huan-ang Gao, Nan Gao, Chun Yu, Yuanchun Shi, and Yuntao Wang. 2025. Computing with Smart Rings: A Systematic Literature Review. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 9, 3, Article 137 (Sept. 2025), 54 pages. doi:10.1145/3749480
@@
- Windlin et al. (2022) Charles Windlin, Kristina Höök, and Jarmo Laaksolahti. 2022. SKETCHING SOMA BITS. In Proceedings of the 2022 ACM Designing Interactive Systems Conference (Virtual Event, Australia) (DIS ’22). Association for Computing Machinery, New York, NY, USA, 1758–1772. doi:10.1145/3532106.3533510
@@
- Yu et al. (2025) Tianhong Catherine Yu, Manru Mary Zhang, Luis Miguel Malenab, Chi-Jung Lee, Jacky Hao Jiang, Ruidong Zhang, François Guimbretière, and Cheng Zhang. 2025. SeamFit: Towards Practical Smart Clothing for Automatic Exercise Logging. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 9, 1, Article 24 (March 2025), 22 pages. doi:10.1145/3712287
@@
- Zhang and Sra (2021) Bowen Zhang and Misha Sra. 2021. PneuMod: A Modular Haptic Device with Localized Pressure and Thermal Feedback. In Proceedings of the 27th ACM Symposium on Virtual Reality Software and Technology (Osaka, Japan) (VRST ’21). Association for Computing Machinery, New York, NY, USA, Article 30, 7 pages. doi:10.1145/3489849.3489857
@@
- Zhou et al. (2025) Ran Zhou, Jianru Ding, Chenfeng Gao, Wanli Qian, Benjamin Erickson, Madeline Balaam, Daniel Leithinger, and Ken Nakagaki. 2025. Shape-Kit: A Design Toolkit for Crafting On-Body Expressive Haptics. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (CHI ’25). Association for Computing Machinery, New York, NY, USA, Article 504, 26 pages. doi:10.1145/3706598.3713981